# Telegrapher’s Generative Model via Kac Flows

### v2

@@ 1 Introduction
Among generative neural models, flow-based techniques as score-based diffusion and flow matching stand out for their simple applicability and good scaling properties. The general idea is to construct a probability flow which starts in a simple latent distribution and leads to a possibly complicated target distribution. A neural network is then trained to learn this flow while using only a few samples from the target. In score-based diffusion models [33, 34], the so-called score is learned which directs the drift-diffusion flow backwards in time from the latent to the target via a certain reverse SDE. In flow matching [17, 18, 19] the velocity field is learned enabling the calculation of the flow’s trajectories via a certain flow ODE. These two approaches produce state-of-the-art results not only in pure sample generation, but can also be incorporated in inverse problem solvers, see, e.g. [22, 44].
@@ 2 Preliminaries: Curves in Wasserstein Spaces
see [1, Definition 1.1.1; Theorem 8.3.1]. If in addition
@@
has a solution $\varphi:I\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\mu_{t}=\varphi(t,\cdot)_{\sharp}\mu_{0}$, see [1, Proposition 8.1.8]. Note that often, absolutely continuous functions in $AC^{1}(I;\mathcal{P}_{2}(\mathbb{R}^{d}))$ with $m\in L_{1}(I)$ in (1) were considered in the literature. However, we need the stronger assumption of square integrability to define appropriate loss functions later.
Given an appropriate velocity field $v_{t}$, we can sample from a target distribution by starting in a sample $x_{0}$ from the latent distribution and applying an ODE solver to (5). However, in this paper, we will reverse the flow direction by starting in the target distribution and ending (approximately) in the latent distribution at some final time $t=T$, similar to diffusion models. If we have access to the forward velocity $v_{t}$, then the reverse flow is simply given by the opposite velocity field $-v_{T-t}$.
@@ 3 Motivation
see [41, Example 4.12], it fulfills the continuity equation (2). Unfortunately, (4) is not fulfilled due to the singularity at $t=1$. Nevertheless, the trajectory $\varphi_{t}$ of the ODE (5) is given by
@@
where $Z_{t}$ is the normalization constant. There are a variety of approaches for approximating the vector field corresponding to the linear interpolation, see, e.g., [24, 28] for kernel-based methods. However, for certain asymmetric constellations of prior and target measure, a mode collapse of the neural model was numerically observed in [25]. Indeed, it was analytically shown in [5] that for the Laplacian $p_{0}\propto e^{-|x|}$ centered at 00 and a Laplacian-like target distribution $p_{1}\propto e^{-2\min\{|x|,|x-m|\}}$ with a mode at 00 and a second one at $m>0$, the velocity norm grows exponentially with $m$, i.e. $\|v_{1}\|_{L_{2}(p_{1})}^{2}\gtrsim e^{\alpha m}$. Using Gaussian mixtures instead of Laplacians, an explosion of the type $\int_{0}^{1}\|v_{t}\|_{L_{2}(p_{t})}^{2}\,\mathrm{d}t\gtrsim m^{4}e^{\alpha m^{2}}$ was proven in [10, Proposition 9].
@@
$-\Delta\varphi(x,z)=p_{\rm data}(x)\delta_{0}(z),\quad x\in\mathbb{R}^{d},z\in\mathbb{R}.$ (17)
@@
where $S_{d}$ is the surface area of the $d$-sphere. For the uniform distribution $p_{\rm data}$ on $[0,1]^{d}$, this results for any $x\in[0,1]^{d}$ and $z=0$ in
@@
Similarly to the diffusion case, the explosion (19) at heights $z$ close to the data ($z=0$) makes suitable truncations necessary when approximating the velocity field by a neural network, see [42, Ch. 3.2].
@@ 4 Telegrapher’s Equation and Kac Model in 1D
Here, $a>0$ is the damping coefficient, and $c>0$ describes the velocity of the wave front. Historically, the equation (4) originates from the modeling of voltage and current in electrical transmission lines (telegraphs). It is known that if the damping $a$ and velocity $c$ go suitably to infinity, see Remark 9, then the diffusion model is retrieved. Hence, diffusion can be seen as "an infinitely $a$-damped wave with infinite propagation speed $c$". The idea of particles traveling with infinite speed has already encountered resistance in the physics community [4, 40, 7, 38], since it violates Einstein’s laws of relativity. But also in the field of generative modeling, this can be seen as a reason for the explosion of velocity fields of diffusion models. Hence, if we restrict ourselves to above model (4), where waves travel only with finite speed $c$, we expect the velocity field $v_{t}$ to behave well. Indeed, we will show that the telegrapher’s model does not admit a velocity explosion, but shares the same long-time asymptotics as diffusion.
@@ 4.1 Kac Flow in 1D
where $\varepsilon_{k}$ are random variables taking the value $-1$ with probability $a\Delta t$ and $1$ with probability $1-a\Delta t$, or alternatively
@@
We refer to $\tau_{t}$ as random time. Finally, for a (symmetrically) random initial direction, the displacement at time $t$ is determined by the Kac process starting in 00,
@@
where ${\rm B}_{\frac{1}{2}}\sim B(1,\frac{1}{2})$ is a Bernoulli random variable taking the values ±$\pm 1$. Naturally, for a random variable $X_{0}$ we call
@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [27] and [23] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [32, Chapter 4.4].
@@
The measure $\mu_{t}$ is exactly the probability distribution of the Kac process $K(t)$ starting in 00 from (25), see [13, Section 2].
Note that the Kac process $K(t)$ starting in 00 does not admit a density itself.
@@
By [13, Section 2], the absolute continuous part of the distribution of the initially right-moving and initially left-moving Kac process ±$\pm S(t)$ starting in 00 is given by
@@
respectively, and it holds $\tilde{u}=\frac{1}{2}(\tilde{u}_{+}+\tilde{u}_{-})$ with $\tilde{u}$ in (33). By [15, Eq. (12) & (11)], the marginal densities $\tilde{u}_{\pm}$ satisfy the PDEs
@@
$\partial_{t}(\tilde{u}_{+}+\tilde{u}_{-})+\,c\,\partial_{x}(\tilde{u}_{+}-\,\tilde{u}_{-})$ $=0.$
@@
$-\int_{-x}^{x}\partial_{t}\tilde{u}(t,y)\,\mathrm{d}y=\int_{-x}^{x}\partial_{x}\tilde{J}(t,y)\,\mathrm{d}y=2\tilde{J}(t,x)\quad\text{ for all }x\in(-ct,ct).$ (48)
@@
Now we can choose a sequence of test functions $\phi_{n}\in C_{c}^{\infty}(\mathbb{R})$ such that $\phi_{n}\equiv 0$ on $[-ct+\frac{1}{n},ct-\frac{1}{n}]$, $\|\phi_{n}\|_{\infty}<\frac{1}{n}$, $\phi_{n}^{\prime}(ct)=1$ and $\phi_{n}^{\prime}(-ct)=-1$. Then the anti-symmetry $v(t,-ct)=-v(t,ct)$ and above identities in the limit $n\to\infty$ yield
@@
and hence, $v(t,ct)=c$. By anti-symmetry, it holds $v(t,-ct)=-c$, and the proof of (43) is finished. ∎
@@ 5.1 Kac’s Insertion Method
This claim was again conjectured in [15] and a proof was given in [9, Theorem 4] using semigroup theory.
@@
In 1D, the solution of (5) is given by $w_{c}(t,x)=\frac{1}{2}(f_{0}(x+ct)+f_{0}(x-ct))$, so that (57) reproduces (27). Unfortunately, in higher dimensions, the telegrapher’s solution (57) might not stay a probability density anymore. For example, the mass $\int_{\mathbb{R}^{3}}u(t,x)\,\mathrm{d}x$ in 3D behaves like $1-e^{-t}$, see [38, Ch. 2.2]. $\diamond$
@@
$h(t,x):=\mathbb{E}\left[w\left(\sigma W(t),x\right)\right],\quad\sigma>0$ (60)
@@
$\partial_{t}h(t,x)$ $=\frac{\sigma^{2}}{2}\Delta h(t,x),$ (61)
@@
The statement can be found in [13], and a proof was given in [9, Theorem 5].
@@
It is well-known that the Laplacian $\Delta:H^{2}(\mathbb{R}^{d})\subset L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ is a self-adjoint linear operator with non-positive spectrum $\sigma(\Delta)=(-\infty,0]$. By [21, Corollary 2] the solution of (5) with $c=1$ is given by $w(t,\cdot)=C(t)f_{0}$, where $(C(t))_{t\geq 0}$ is a uniformly bounded so-called cosine operator function with $\|C(t)\|_{L_{2}\to L_{2}}\leq 1$ for all $t\geq 0$. In particular, it follows that $\|w(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq\|f_{0}\|_{L_{2}(\mathbb{R}^{d})}$ is uniformly bounded in $t$. Now, [13, Theorem 2] yields the claim. ∎
@@ 5.2 Multi-Dimensional Kac Process
For a random variable $\mathbf{X}_{0}\in\mathbb{R}^{d}$, $d\geq 1$, we consider i.i.d. [^2] We use i.i.d. Kac processes just for simplicity. But our framework also allows for independent processes $K_{i}$ each with their own parameters $(a_{i},c_{i})$. copies $K^{i}(t)$, $i=1,...,d$, of a Kac process (25) starting in $\mathbf{0}$. Then, we call
@@
ii) By definition (1) and Part i), we see immediately that $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve. Then it follows from [1, Theorem 8.3.1] that there is a (optimal) velocity field in the continuity equation fulfilling (65).
[…]
@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
By Proposition 10, $\mu_{t}(\cdot|x_{0})$ admits a velocity field $v_{t}(\cdot|x_{0})$ satisfying the continuity equation (2). By construction, the components $(X_{t}^{1},...,X_{t}^{d})$ of $\mathbf{X}_{t}$ are mutually independent. Hence, $\mu_{t}(x|x_{0})=\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}|x_{0}^{i})$, where the one-dimensional probability flows $\mu_{t}^{i}(\cdot|x_{0}^{i})$ fulfill a continuity equation with 1D velocity fields $v_{t}^{i}(\cdot|x_{0}^{i})$. By Lemma 12, it holds the decomposition (86), and Theorem 4 immediately gives the formula (91). Finally, [19, Theorem 3.3], resp. [41, Theorem 6.3] provides (92). ∎
@@
Having identified the conditional vector field $v_{t}(\cdot|x_{0})$ for the conditional distribution $\mu_{t}(\cdot|x_{0})$ of the Kac process, we get the following result by employing [17], see also [41, Theorem 5.1, Remark 5.3], with slight adjustments to our setting.
@@
where $\mathcal{U}(0,T)\times_{t}u_{t}$ is understood in the sense of disintegration, see [41, Ch. 2.3], but here we have no direct access to $v$. However, following the same lines as in [17], we see that
@@ 7 Numerical Experiments
In this section, we use the flow matching framework to train a neural network to approximate the velocity field from Corollary 14 by minimizing the CFM loss given in (95). To this end, we compute the conditional velocity $v_{t}(x|x_{0})$ via its analytic formula (91). In order to sample from $\mathbf{X}_{t}\sim u_{t}(\cdot|x_{0})$, where $x_{0}$ is fixed, we only need to sample i.i.d. components $X_{t}^{i}$ of the Kac process starting in 00. This can be done efficiently, since we have access to the analytic form (32) of the one-dimensional Kac distribution starting in 00 (use e.g. inverse transform sampling). Further, if we wish to sample fixed realizations of a Kac process for varying times $t$, we can efficiently simulate the random walk itself as explained in Appendix A.1.
@@
We trained each model for $500$k iterations, and used the negative log-likelihood as a validation loss. In Figure 5 we plot the resulting generated samples and report the associated training iterations. The complete experimental setup can be found in the Appendix A.2. For the backward evolution of the learned $(25,5)$-Kac flow, see Figure 4.
@@
In contrast to our Kac generative model, the diffusion based model is not able to precisely approximate the Dirac-like modes, even with long training times and vanishing truncation times [^4] By the explosion (12) time truncations were necessary, since otherwise the associated adaptive step size ODE solver diverged with intended step size going to 00. $\epsilon\approx 0$. Clearly, high choices of parameters $(a,c)$ lead to diffusion-like results as theoretically underlined in Theorem 11.

### v3

@@ 1 Introduction
Among generative neural models, flow-based techniques as score-based diffusion and flow matching stand out for their simple applicability and good scaling properties. The general idea is to construct a probability flow which starts in a simple latent distribution and leads to a possibly complicated target distribution. A neural network is then trained to learn this flow while using only a few samples from the target. In score-based diffusion models [35, 36], the so-called score is learned which directs the drift-diffusion flow backwards in time from the latent to the target via a certain reverse SDE. In flow matching [19, 20, 21] the velocity field is learned enabling the calculation of the flow’s trajectories via a certain flow ODE. These two approaches produce state-of-the-art results not only in pure sample generation, but can also be incorporated into inverse problem solvers, see e.g. [24, 46]. A unifying perspective can be given via the ”stochastic interpolants” framework, see [1, 2].
@@ 2 Preliminaries: Curves in Wasserstein Spaces
see [3, Definition 1.1.1; Theorem 8.3.1]. If in addition
@@
has a solution $\varphi:I\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\mu_{t}=\varphi(t,\cdot)_{\sharp}\mu_{0}$, see [3, Proposition 8.1.8]. Note that often, absolutely continuous functions in $AC^{1}(I;\mathcal{P}_{2}(\mathbb{R}^{d}))$ with $m\in L_{1}(I)$ in (1) were considered in the literature. However, we need the stronger assumption of square integrability to define appropriate loss functions later.
Given an appropriate velocity field $v_{t}$, we can sample from a target distribution by starting in a sample $x_{0}$ from the latent distribution and applying an ODE solver to (5). However, in this paper, we will reverse the flow direction by starting in the target distribution and ending (approximately) in the latent distribution at some final time $t=T$, similar to diffusion models. If we have access to the forward velocity $v_{t}$, then the reverse flow is simply given by the opposite velocity field $v_{T-t}$.
@@ 3 Motivation
see [43, Example 4.12], it fulfills the continuity equation (2). Unfortunately, (4) is not fulfilled due to the singularity at $t=1$. Nevertheless, the trajectory $\varphi_{t}$ of the ODE (5) is given by
@@
where $Z_{t}$ is the normalization constant. There are a variety of approaches for approximating the vector field corresponding to the linear interpolation, see, e.g., [26, 30] for kernel-based methods. However, for certain asymmetric constellations of prior and target measure, a mode collapse of the neural model was numerically observed in [27]. Indeed, it was analytically shown in [7] that for the Laplacian $p_{0}\propto e^{-|x|}$ centered at 0 and a Laplacian-like target distribution $p_{1}\propto e^{-2\min\{|x|,|x-m|\}}$ with a mode at 0 and a second one at $m>0$, the velocity norm grows exponentially with $m$, i.e. $\|v_{1}\|_{L_{2}(p_{1})}^{2}\gtrsim e^{\alpha m}$. Using Gaussian mixtures instead of Laplacians, an explosion of the type $\int_{0}^{1}\|v_{t}\|_{L_{2}(p_{t})}^{2}\,\mathrm{d}t\gtrsim m^{4}e^{\alpha m^{2}}$ was proven in [12, Proposition 9].
@@
$\Delta\varphi(x,z)=p_{\rm data}(x)\delta_{0}(z),\quad x\in\mathbb{R}^{d},z\in\mathbb{R}.$ (17)
@@
where $S_{d}$ is the surface area of the $d$-sphere. For the uniform distribution $p_{\rm data}$ on $0,1]^{d}$, this results for any $x\in[0,1]^{d}$ and $z=0$ in
@@
Similarly to the diffusion case, the explosion (19) at heights $z$ close to the data ($z=0$) makes suitable truncations necessary when approximating the velocity field by a neural network, see [44, Ch. 3.2].
@@ 4 Telegrapher’s Equation and Kac Model in 1D
Here, $a>0$ is the damping coefficient, and $c>0$ describes the velocity of the wave front. Historically, the equation (4) originates from the modeling of voltage and current in electrical transmission lines (telegraphs). It is known that if the damping $a$ and velocity $c$ go suitably to infinity, see Remark 9, then the diffusion model is retrieved. Hence, diffusion can be seen as "an infinitely $a$-damped wave with infinite propagation speed cc". The idea of particles traveling with infinite speed has already encountered resistance in the physics community [6, 42, 9, 40], since it violates Einstein’s laws of relativity. But also in the field of generative modeling, this can be seen as a reason for the explosion of velocity fields of diffusion models. Hence, if we restrict ourselves to above model (4), where waves travel only with finite speed $c$, we expect the velocity field $v_{t}$ to behave well. Indeed, we will show that the telegrapher’s model does not admit a velocity explosion, but shares the same long-time asymptotics as diffusion.
@@ 4.1 Kac Flow in 1D
where $\varepsilon_{k}$ are random variables taking the value 1-1 with probability $a\Delta t$ and $1$ with probability $1-a\Delta t$, or alternatively
@@
We refer to $\tau_{t}$ as random time. Finally, for a (symmetrically) random initial direction, the displacement at time $t$ is determined by the Kac process starting in 0,
@@
where ${\rm B}_{\frac{1}{2}}\sim B(1,\frac{1}{2})$ is a Bernoulli random variable taking the values $\pm 1$. Naturally, for a random variable $X_{0}$ we call
@@ 4.2 Relation of Kac’ Flow to 1D Telegrapher’s Equation
By the next lemma, there is an analytic representation of the distributional solution of the telegrapher’s equation (4). We refer to [29] and [25] for the proof via distributions. For a characterization of measures as (tempered) distributions, see also [34, Chapter 4.4].
@@
The measure $\mu_{t}$ is exactly the probability distribution of the Kac process $K(t)$ starting in 0 from (25), see [15, Section 2].
Note that the Kac process $K(t)$ starting in 0 does not admit a density itself.
@@
By [15, Section 2], the absolute continuous part of the distribution of the initially right-moving and initially left-moving Kac process $\pm S(t)$ starting in 0 is given by
@@
[…]
@@
[…]
@@
[…]
@@
Now we can choose a sequence of test functions $\phi_{n}\in C_{c}^{\infty}(\mathbb{R})$ such that $\phi_{n}\equiv 0$ on $ct+\frac{1}{n},ct-\frac{1}{n}]$, $\|\phi_{n}\|_{\infty}<\frac{1}{n}$, $\phi_{n}^{\prime}(ct)=1$ and $\phi_{n}^{\prime}(-ct)=-1$. Then the anti-symmetry v(t,ct)=v(t,ct)v(t,-ct)=-v(t,ct) and above identities in the limit $n\to\infty$ yield
@@
and hence, $v(t,ct)=c$. By anti-symmetry, it holds v(t,ct)=cv(t,-ct)=-c, and the proof of (43) is finished. ∎
@@ 5.1 Kac’s Insertion Method
This claim was again conjectured in [17] and a proof was given in [11, Theorem 4] using semigroup theory.
@@
In 1D, the solution of (5) is given by $w_{c}(t,x)=\frac{1}{2}(f_{0}(x+ct)+f_{0}(x-ct))$, so that (57) reproduces (27). Unfortunately, in higher dimensions, the telegrapher’s solution (57) might not stay a probability density anymore. For example, the mass $\int_{\mathbb{R}^{3}}u(t,x)\,\mathrm{d}x$ in 3D behaves like $1-e^{-t}$, see [40, Ch. 2.2]. $\diamond$
@@
$h(t,x):=\mathbb{E}\left[w\left(\sigma W(t),x\right)\right],\quad\sigma>0,$ (60)
@@
$\partial_{t}h(t,x)$ $=\frac{\sigma^{2}}{2}\Delta h(t,x),\quad x\in\mathbb{R}^{d},~t>0,$ (61)
@@
The statement can be found in [15], and a proof was given in [11, Theorem 5].
@@
It is well-known that the Laplacian $\Delta:H^{2}(\mathbb{R}^{d})\subset L_{2}(\mathbb{R}^{d})\to L_{2}(\mathbb{R}^{d})$ is a self-adjoint linear operator with non-positive spectrum $\sigma(\Delta)=(-\infty,0]$. By [23, Corollary 2] the solution of (5) with $c=1$ is given by $w(t,\cdot)=C(t)f_{0}$, where $(C(t))_{t\geq 0}$ is a uniformly bounded so-called cosine operator function with $\|C(t)\|_{L_{2}\to L_{2}}\leq 1$ for all $t\geq 0$. In particular, it follows that $\|w(t,\cdot)\|_{L_{2}(\mathbb{R}^{d})}\leq\|f_{0}\|_{L_{2}(\mathbb{R}^{d})}$ is uniformly bounded in $t$. Now, [15, Theorem 2] yields the claim. ∎
@@ 5.2 Multi-Dimensional Kac Process
For a random variable $\mathbf{X}_{0}\in\mathbb{R}^{d}$, $d\geq 1$, we consider i.i.d. [^2] We use i.i.d. Kac processes just for simplicity. But our framework also allows for independent processes $K_{i}$ each with their own parameters $(a_{i},c_{i})$. copies $K^{i}(t)$, i=1,,di=1,...,d, of a Kac process (25) starting in $\mathbf{0}$. Then, we call
@@
ii) By definition (1) and Part i), we see immediately that $(\mu_{t})_{t\geq 0}$ is an absolutely continuous curve. Then it follows from [3, Theorem 8.3.1] that there is a (optimal) velocity field in the continuity equation fulfilling (65).
iii) Let $\bar{\Omega}_{0}\coloneqq\mathrm{supp}(P_{\mathbf{X}_{0}})$. Consider the covering $\bar{\Omega}_{t}\coloneqq\bigcup_{x\in\bar{\Omega}_{0}}\mathbb{B}_{d^{\frac{1}{2}}ct}(x)$ of $\bar{\Omega}_{0}$ with balls of radius $d^{\frac{1}{2}}ct$. By (5.2) it follows that $\mathbf{X}_{t}\in\bar{\Omega}_{t}$ $\mathbb{P}$-almost surely, yielding the claim $\mathrm{supp}(P_{\mathbf{X}_{t}})\subseteq\bar{\Omega}_{t}$. ∎
@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
By Proposition 10, $\mu_{t}(\cdot|x_{0})$ admits a velocity field $v_{t}(\cdot|x_{0})$ satisfying the continuity equation (2). By construction, the components $(X_{t}^{1},...,X_{t}^{d})$ of $\mathbf{X}_{t}$ are mutually independent. Hence, $\mu_{t}(x|x_{0})=\prod_{i=1}^{d}\mu_{t}^{i}(x^{i}|x_{0}^{i})$, where the one-dimensional probability flows $\mu_{t}^{i}(\cdot|x_{0}^{i})$ fulfill a continuity equation with 1D velocity fields $v_{t}^{i}(\cdot|x_{0}^{i})$. By Lemma 12, it holds the decomposition (86), and Theorem 4 immediately gives the formula (91). Finally, [21, Theorem 3.3], resp. [43, Theorem 6.3] provides (92). ∎
@@
Having identified the conditional vector field $v_{t}(\cdot|x_{0})$ for the conditional distribution $\mu_{t}(\cdot|x_{0})$ of the Kac process, we get the following result by employing [19], see also [43, Theorem 5.1, Remark 5.3], with slight adjustments to our setting.
@@
where $\mathcal{U}(0,T)\times_{t}u_{t}$ is understood in the sense of disintegration, see [43, Ch. 2.3], but here we have no direct access to $v$. However, following the same lines as in [19], we see that
@@ 7 Numerical Experiments
In this section, we use the flow matching framework to train a neural network to approximate the velocity field from Corollary 14 by minimizing the CFM loss given in (95). To this end, we compute the conditional velocity $v_{t}(x|x_{0})$ via its analytic formula (91). In order to sample from $\mathbf{X}_{t}\sim u_{t}(\cdot|x_{0})$, where $x_{0}$ is fixed, we only need to sample i.i.d. components $X_{t}^{i}$ of the Kac process starting in 0. This can be done efficiently, since we have access to the analytic form (32) of the one-dimensional Kac distribution starting in 0 (use e.g. inverse transform sampling). Further, if we wish to sample fixed realizations of a Kac process for varying times $t$, we can efficiently simulate the random walk itself as explained in Appendix A.1.
@@
We trained each model for 500500k iterations, and used the negative log-likelihood as a validation loss. In Figure 5 we plot the resulting generated samples and report the associated training iterations. The complete experimental setup can be found in the Appendix A.2. For the backward evolution of the learned $(25,5)$-Kac flow, see Figure 4.
@@
In contrast to our Kac generative model, the diffusion based model is not able to precisely approximate the Dirac-like modes, even with long training times and vanishing truncation times [^4] By the explosion (12) time truncations were necessary, since otherwise the associated adaptive step size ODE solver diverged with intended step size going to 0. $\epsilon\approx 0$. Clearly, high choices of parameters $(a,c)$ lead to diffusion-like results as theoretically underlined in Theorem 11.